OpenStudy (anonymous):

draw a slope using the equation -3/4 passing through the point (-5,1)

OpenStudy (hanner_b_nanner):

I'm pretty sure you can just place a point on (-5,1), and then use rise/run=-3/4 to finish making the line! :)

OpenStudy (the_fizicx99):

An equation? Hm, do you mean to say the slope is -3/4? and Use the points (-5,1) y-y1 = m(x-x1)

OpenStudy (anonymous):

Well if the slope is -3/4 and you have to find the equation then: y = mx + b y = -3/4x + b 1 = -3/4(-5) + b 1 = 3 3/4 + b -2 3/4 = b y = -3/4x -2 3/4 i think :)